OpenStudy (anonymous):

can someone walk me through this please? use a limit and evaluate the integral by hand to prove that the following:

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}x^{-1.5}dx=2\]

OpenStudy (anonymous):

@LolWolf and @mathslover maybe since you've helped me before?

OpenStudy (anonymous):

By hand...? I don't quite get what that means, but: \[ \lim_{a\to \infty}\left(\int_1^ax^{-1.5}\;dx\right)=\lim_{a\to \infty}\left[-\frac{x^{-.5}}{.5}\right]_1^a \]Try evaluating that.

OpenStudy (anonymous):

I think it just means without using the calculator to solve

OpenStudy (anonymous):

Sorry I'm lost